OpenStudy (anonymous):

Correct answer for a medal. What is the third term in the binomial expansion of (3x + y3)4? 54x^2*y^3 18x^2*y^3 18x^2*y^6 54x^2*y^6

OpenStudy (anonymous):

I keep getting getting 18x^2*y^6 but it says that my answer is wrong.

OpenStudy (fibonaccichick666):

can you show me your work?

OpenStudy (fibonaccichick666):

I can't fix it without seeing where exactly you are going wrong

OpenStudy (anonymous):

I think I have been writing the formula incorrectly. :/

OpenStudy (fibonaccichick666):

ok well, tell me what you got. I need to see it, unfortunately I am not a telepath

OpenStudy (anonymous):

@FibonacciChick666 4C2 a^2 b^3 but I don't know how to solve for C Sorry for the wait.

OpenStudy (fibonaccichick666):

ohk yea wrong formula, try this: \[(x+y)^4 \;=\; x^4 \,+\, 4 x^3y \,+\, 6 x^2 y^2 \,+\, 4 x y^3 \,+\, y^4.\]

OpenStudy (anonymous):

The 4 and the 2 next to the C are supposed to be subscripts.

OpenStudy (fibonaccichick666):

that is the binomial theorem

OpenStudy (fibonaccichick666):

so which is the third term?

OpenStudy (anonymous):

6x^2 y^2

OpenStudy (fibonaccichick666):

good, so now the 'x' and 'y' are actually loose terms here they mean whatever is in the first slot and whatever is in the second slot. so what are our 'x' and 'y'?

OpenStudy (anonymous):

2 and 2?

OpenStudy (fibonaccichick666):

where did you get the 2's?

OpenStudy (anonymous):

x^2 and y^2 from the third term

OpenStudy (fibonaccichick666):

ok so let me rephrase, instead of \[(x+y)^4 \;=\; x^4 \,+\, 4 x^3y \,+\, 6 x^2 y^2 \,+\, 4 x y^3 \,+\, y^4.\] let's say\[(a+b)^4 \;=\; a^4 \,+\, 4 a^3b \,+\, 6 a^2 b^2 \,+\, 4 a b^3 \,+\, b^4.\]

OpenStudy (fibonaccichick666):

so we have \[(3x+y^3)^4\] What is our a, and what is our b?

OpenStudy (anonymous):

I'm sorry, but I am confused... 1 and 3?

OpenStudy (fibonaccichick666):

it's ok so let's look at this one (5x+4y) in this case our a=5x and our b=4y

OpenStudy (fibonaccichick666):

do you understand what I mean now?

OpenStudy (anonymous):

oh ok so a=3x and b=y^3 right?

OpenStudy (fibonaccichick666):

yep now just plug it in

OpenStudy (anonymous):

6 * 3x^2 * y^6

OpenStudy (fibonaccichick666):

not quite remember parentheses

OpenStudy (anonymous):

6*3x^4*y^8